OpenStudy (anonymous):

What is the solution of the equation? 6x – 3 = 5x – 5 A. –4 B. –2 C. 0 D. –1

OpenStudy (anonymous):

this idk

OpenStudy (anonymous):

a i think

OpenStudy (anonymous):

B

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

b then

OpenStudy (australopithecus):

6x - 3 = 5x - 5 6x - 3 - 6x = 5x - 5 - 6x - 3 + 0 = -1x - 5 -3 + 5 = -x - 5 + 5 2 = -x + 0 2 = -x 2*-1 = -x(-1) -2 = x

OpenStudy (australopithecus):

remember what you do to one side of the equation you must to do to the other to maintain the equality

OpenStudy (anonymous):

6x-3=5x-5 +3 +3 6x=5x-2 -5x=-5x 1x=-2 /1 /1 -2/1=-2

OpenStudy (australopithecus):

and also remember a -*- = +